# Non-Convex Self-Concordant Functions: Practical Algorithms and Complexity Analysis

### v1

@@ 2.1 Examples
Many problems admit loss functions that can be modified to be $F$-based self-concordant for some function $F$. For our next example, we define $h_{\alpha}(x)=\alpha^{-1}(\log(1+e^{\alpha x})+\log(1+e^{-\alpha x}))$. For small $\alpha>0$, $h_{\alpha}$ serves as a smooth approximation to $|x|$ [Schmidt-2007-Fast], and it is $0.5\alpha$-self-concordant. After some calculation, it can be verified that $\frac{\partial^{2}}{\partial x^{2}}h_{\alpha}(x)=\frac{\alpha}{2}(1-\tanh^{2}(\frac{\alpha x}{2}))$, $\frac{\partial^{3}}{\partial x^{3}}h_{\alpha}(x)=-\frac{\alpha^{2}\tanh(\alpha x/2)}{2\cosh^{2}(\alpha x/2)}$, and that $|\frac{\partial^{3}}{\partial x^{3}}h_{\alpha}(x)|\leq\alpha(\frac{\partial^{2}}{\partial x^{2}}h_{\alpha}(x))$ for $x\geq 0$.
Sparse dictionary learning
As introduced in [Mairal-2008-Supervised], we consider the optimization problem in the following form of
$\min_{\|d_{i}\|\leq 1}f(D,r_{1},\ldots,r_{K})=\sum_{i=1}^{K}\|x_{i}-Dr_{i}\|^{2}+\lambda\left(\sum_{i=1}^{K}\sum_{j=1}^{n}|r_{i}^{(j)}|\right),$
@@
where $x_{i}\in\mathbb{R}^{m}$, $r_{i}\in\mathbb{R}^{n}$, $D\in\mathbb{R}^{m\times n}$, $d_{i}$ denotes the $i$-th column of $D$, $r_{i}^{(j)}$ denotes the $j$-th entry of $r_{i}$, and $\lambda>0$. We apply $h_{\alpha}$ to smooth the absolute value function and the self-concordant function $g(d_{i})=-\log(1-\|d_{i}\|^{2})$ (see [Nesterov-2018-Lectures, Theorem 5.1.4]) to impose $\|d_{i}\|<1$. This yields the approximation in the following form of
$f_{\alpha,\mu}(D,r_{1},\ldots,r_{K})=\sum_{i=1}^{K}\|x_{i}-Dr_{i}\|_{2}^{2}+\lambda\left(\sum_{i=1}^{K}\sum_{j=1}^{n}h_{\alpha}(r_{i}^{(j)})\right)+\mu\left(\sum_{i=1}^{n}g(d_{i})\right).$
Propositions 2.4 and 2.5 imply that $f_{\alpha,\mu}$ is $F$-based 11-self-concordant.
Nonnegative matrix factorization
Let $\mathbb{R}_{+}=\{x\in\mathbb{R}:x\geq 0\}$, $\mathbb{R}_{++}=\{x\in\mathbb{R}:x>0\}$. We show that two NMF formulations [Lee-2000-Algorithms] using Frobenius norm in Eq. (2) and Kullback–Leibler (KL) divergence in Eq. (3), both fit into our nonconvex self-concordant framework. Indeed, we have
$\min_{X\in\mathbb{R}_{++}^{m\times r},Y\in\mathbb{R}_{++}^{r\times n}}f_{1}(X,Y)=\tfrac{1}{2mn}\|Z-XY\|_{2}^{2}=\tfrac{1}{2mn}\left(\sum_{i=1}^{m}\sum_{j=1}^{n}\left(Z_{ij}-\sum_{k=1}^{r}X_{ik}Y_{kj}\right)^{2}\right),$ (2)
@@
where $Z\in\mathbb{R}^{m\times n}_{+}$. By Proposition 2.5, $f_{1}$ is $\ell F$-based $1$-self-concordant for $F(X,Y)=(\|X\|_{F}^{2}+\|Y\|_{F}^{2}+1)^{2}-\sum_{i,k}\log(X_{ik})-\sum_{k,j}\log(Y_{kj})$ and some $\ell>0$. Another formulation using KL divergence is in the following form of
$\min_{X\in\mathbb{R}_{++}^{m\times r},Y\in\mathbb{R}_{++}^{r\times n}}f_{2}(X,Y)=\tfrac{1}{mn}\left(\sum_{i=1}^{m}\sum_{j=1}^{n}D\left(Z_{ij}\Big\|\sum_{k=1}^{r}X_{ik}Y_{kj}\right)\right),$ (3)
where $D(x\|y)=x\log(x/y)-x+y$.
@@
We prove that $f_{2}(X,Y)$ is $F$-based 11-self-concordant for some convex function $F$ on $\mathbb{R}_{++}^{m\times r}\times\mathbb{R}_{++}^{r\times n}$. Indeed, we let $g(x,y)=-\log(x^{\top}y)$ and $G(x,y)=-\sum_{i=1}^{r}\log(x_{i}y_{i})$. By Proposition 2.4, it suffices to prove that $g$ is $\tau G$-based $1$-self-concordant for some $\tau>3$. Fixing $(x,y)\in\mathbb{R}_{++}^{r}\times\mathbb{R}_{++}^{r}$ and $h=(h_{x},h_{y})\in\mathbb{R}^{r}\times\mathbb{R}^{r}$, we define
$s=x^{\top}y,\quad L=h_{x}^{\top}y+h_{y}^{\top}x,\quad u_{i}=\tfrac{|(h_{x})_{i}|}{x_{i}},\quad v_{i}=\tfrac{|(h_{y})_{i}|}{y_{i}},\quad w_{i}=\tfrac{x_{i}y_{i}}{s}.$
@@
$\tfrac{|L|}{s}\leq\sum_{i=1}^{r}w_{i}(|u_{i}|+|v_{i}|)\leq\sqrt{2\left(\sum_{i=1}^{r}w_{i}(u_{i}^{2}+v_{i}^{2})\right)}\leq\sqrt{2\left(\sum_{i=1}^{r}(u_{i}^{2}+v_{i}^{2})\right)},$
and
$\tfrac{|h_{x}^{\top}h_{y}|}{s}\leq\sum_{i=1}^{r}\tfrac{|(h_{x})_{i}||(h_{y})_{i}|}{x_{i}y_{i}}\leq\tfrac{1}{2}\left(\sum_{i=1}^{r}(u_{i}^{2}+v_{i}^{2})\right).$
Thus, we have
$|\nabla^{2}g(x,y)[h,h]|=\left|\tfrac{L^{2}}{s^{2}}-\tfrac{2h_{x}^{\top}h_{y}}{s}\right|\leq 3\left(\sum_{i=1}^{r}(u_{i}^{2}+v_{i}^{2})\right),$
and
$|\nabla^{3}g(x,y)[h,h,h]|=\left|-2\tfrac{L^{3}}{s^{3}}+\tfrac{6Lh_{x}^{\top}h_{y}}{s^{2}}\right|\leq 7\sqrt{2}\left(\sum_{i=1}^{r}(u_{i}^{2}+v_{i}^{2})\right)^{3/2}.$
In addition, we obtain that $\nabla^{2}G(x,y)[h,h]=\sum_{i=1}^{r}(u_{i}^{2}+v_{i}^{2})$ and $|\nabla^{3}G(x,y)[h,h,h]|\leq 2(\sum_{i=1}^{r}(u_{i}^{3}+v_{i}^{3}))\leq 2(\sum_{i=1}^{r}(u_{i}^{2}+v_{i}^{2}))^{3/2}$. Thus, we have
$\nabla^{2}(g+\tau G)[h,h]$ $\geq$ $(\tau-3)\left(\sum_{i=1}^{r}(u_{i}^{2}+v_{i}^{2})\right),$
$|\nabla^{3}(g+\tau G)[h,h,h]|$ $\leq$ $(2\tau+7\sqrt{2})\left(\sum_{i=1}^{r}(u_{i}^{2}+v_{i}^{2})\right)^{3/2}.$
This implies that $|\nabla^{3}(g+\tau G)[h,h,h]|\leq 2(\nabla^{2}(g+\tau G)[h,h])^{3/2}$ for some $\tau>3$. This yields the desired result.
2.2 Descent Inequality
We establish the descent inequality for $F$-based $\kappa$-self-concordant functions where $F$ is a convex function. For simplicity, we define two auxiliary functions $\omega(z)$ and $\omega_{\star}(z)$ on $\mathbb{R}_{+}$ as follows,
$\omega(z)=z-\log(1+z)\textnormal{ and }\omega_{\star}(z)=\begin{cases}-z-\log(1-z)&\textnormal{if }z<1,\\ +\infty&\textnormal{otherwise}.\end{cases}$
In addition, suppose that $f$ is lower bounded, we define
@@ 2.2 Descent Inequality
Suppose that $f$ is $F$-based $\kappa$-self-concordant. Then, for any $x\in\textbf{dom}(f)$ and $d\in\mathbb{R}^{n}$ satisfying $\nabla f(x)^{\top}d\leq 0$, we define
$\rho=-\nabla f(x)^{\top}d,\quad\delta=\nabla^{2}f(x)[d,d],\quad\Delta=\nabla^{2}F(x)[d,d],\quad\eta=\tfrac{\rho}{\sqrt{\Delta+\delta}}.$
For any $t\in[0,\frac{1}{\kappa\sqrt{\delta+\Delta}})$, we have $x+td\in\textbf{dom}(f)$ and
If $F$ is convex, we have
@@
Suppose that $\bar{t}=\frac{\rho}{\delta+\Delta+\kappa\rho\sqrt{\delta+\Delta}}<\frac{1}{\kappa\sqrt{\delta+\Delta}}$ minimizes the right-hand side of Eq. (5). Then, we have $f(x+td)\leq f(x)$ for all $t\in[0,\bar{t}]$ and
$f(x+\bar{t}d)\leq f(x)-\rho\bar{t}+\kappa^{-2}\omega_{\star}(\kappa\bar{t}\sqrt{\delta+\Delta})=f(x)-\kappa^{-2}\omega(\kappa\eta).$ (6)
If $f$ is lower bounded, we have $\kappa^{-2}\omega(\kappa\eta)\geq\frac{\eta^{2}}{2(1+\Gamma_{f}(x))}$. If $F$ is convex and $\kappa_{F}$-self-concordant, we have
@@
It follows from [Nesterov-2018-Lectures, Theorems 5.1.5 and 5.1.9] that $x+td\in\textbf{dom}(f)$ and
@@
Since $F$ is convex, we have $F(x)+t\nabla F(x)^{\top}d-F(x+td)\leq 0$. Plugging this into the above inequality yields Eq. (5). We define
@@
$g^{\prime}(t)=-\rho+\left(\tfrac{t(\delta+\Delta)}{1-\kappa t\sqrt{\delta+\Delta})}\right).$
Thus, $g^{\prime}(t)\leq 0$ for all $t\in[0,\bar{t}]$. Since $g(0)=f(x)$, we have $f(x+td)\leq g(t)\leq f(x)$ for all $t\in[0,\bar{t}]$. Plugging $\bar{t}$ into (5) yields Eq. (6). In addition, Eq. (6) implies
@@
Thus, $\kappa\eta\leq\Gamma_{f}(x)$. Since $w(z)\geq\frac{z^{2}}{2(1+\Gamma_{f}(x))}$ for $0\leq z\leq\Gamma_{f}(x)$, we have $\kappa^{-2}\omega(\kappa\eta)\geq\frac{\eta^{2}}{2(1+\Gamma_{f}(x))}$. It follows from [Nesterov-2018-Lectures, Theorem 5.1.8] that
@@
which implies Eq. (7). ∎
@@ 3 Algorithms
We develop two algorithms for solving the problem in Eq. (1), where the objective function $f$ is $F$-based $\kappa$-self-concordant and bounded below. Compared to existing methods [Nesterov-2006-Cubic, Cartis-2011-Adaptive, Cartis-2011-Adaptive2, Gratton-2025-Yet], our two algorithms rely on neither the Lipschitz Hessian condition nor the unconstrained domain (i.e., $\textbf{dom}(f)=\mathbb{R}^{n}$). Throughout this paper, we make Assumption 3.1 and impose additional assumptions (e.g., $F$ is self-concordant or quadratic) if necessary.
@@
The function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is bounded below and is $F$-based $\kappa$-self-concordant with respect to a convex function $F:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ for some $\kappa\geq 0$. In addition, $\nabla^{2}(f+F)(x)$ is positive definite for any $x\in\textbf{dom}(f)$.

### v2

@@ 2.1 Examples
Many problems other than POPs involve loss functions that can be modified to be $F$-based self-concordant for some function $F$. Our next example SDL is such a problem. We define $h_{\alpha}(x)=\alpha^{-1}(\log(1+e^{\alpha x})+\log(1+e^{-\alpha x}))$. For large $\alpha>0$, $h_{\alpha}$ serves as a smooth approximation to $|x|$ [68], and it is weakly self-concordant. After some calculation, it can be verified that $\frac{\partial^{2}}{\partial x^{2}}h_{\alpha}(x)=\frac{\alpha}{2}(1-\tanh^{2}(\frac{\alpha x}{2}))$, $\frac{\partial^{3}}{\partial x^{3}}h_{\alpha}(x)=-\frac{\alpha^{2}\tanh(\alpha x/2)}{2\cosh^{2}(\alpha x/2)}$, and that both of $\left|\frac{\partial^{2}}{\partial x^{2}}h_{\alpha}(x)\right|$ and $\left|\frac{\partial^{3}}{\partial x^{3}}h_{\alpha}(x)\right|$ are bounded.
Sparse Dictionary Learning (SDL)
As introduced in [45], we consider the optimization problem
$\min_{\|d_{i}\|\leq 1,i\in N}f(D,r_{1},\ldots,r_{K})=\sum_{i=1}^{K}\|x_{i}-Dr_{i}\|^{2}+\lambda\left(\sum_{i=1}^{K}\sum_{j=1}^{n}|r_{i}^{(j)}|\right),$
@@
where $N=\{1,\ldots,n\}$, $x_{i}\in\mathbb{R}^{m}$, $r_{i}\in\mathbb{R}^{n}$, $D\in\mathbb{R}^{m\times n}$, $d_{i}$ denotes the $i$-th column of $D$, $r_{i}^{(j)}$ denotes the $j$-th entry of $r_{i}$, and $\lambda>0$. We modify $f$ in two ways; specifically, we apply $h_{\alpha}$ to smoothly approximate $|r_{i}^{(j)}|$ and introduce the self-concordant function $g(d_{i})=-\log(1-\|d_{i}\|^{2})$ (see [58, Theorem 5.1.4]) to impose $\|d_{i}\|<1$. This yields the following approximation
$f_{\alpha,\mu}(D,r_{1},\ldots,r_{K})=\sum_{i=1}^{K}\|x_{i}-Dr_{i}\|^{2}+\lambda\left(\sum_{i=1}^{K}\sum_{j=1}^{n}h_{\alpha}(r_{i}^{(j)})\right)+\mu\left(\sum_{i=1}^{n}g(d_{i})\right),$
to $f$. Propositions 2.4 and 2.5 imply that $f_{\alpha,\mu}$ is $F$-based 11-self-concordant.
Nonnegative Matrix Factorization (NMF)
Let $\mathbb{R}_{+}=\{x\in\mathbb{R}:x\geq 0\}$ and $\mathbb{R}_{++}=\{x\in\mathbb{R}:x>0\}$. We consider here two alternative formulations of the problem of approximating a nonnegative matrix $Z$ by the product $XY$ of two nonnegative matrices $X$ and $Y$ [39]. The first formulation measures the difference $D$ between $Z$ and $XY$ by the square of the Frobenius norm of $D$ (i.e., the mean squared error (MSE)), while the second formulation measures $D$ by the Kullback–Leibler (KL) divergence between $Z$ and $XY$; see (2) and (3) below. Both fit into our non-convex self-concordant framework. Indeed, the first formulation yields the optimization problem
$\min_{X\in\mathbb{R}_{++}^{m\times r},Y\in\mathbb{R}_{++}^{r\times n}}f_{\text{MSE}}(X,Y)=\tfrac{1}{2mn}\|Z-XY\|_{F}^{2}=\tfrac{1}{2mn}\left(\sum_{i=1}^{m}\sum_{j=1}^{n}\left(Z_{ij}-\sum_{k=1}^{r}X_{ik}Y_{kj}\right)^{2}\right),$ (2)
@@
where $Z\in\mathbb{R}^{m\times n}_{+}$. By Proposition 2.5, $f_{\text{MSE}}$ is $\ell F_{\text{MSE}}$-based $\kappa$-self-concordant for $F_{\text{MSE}}(X,Y)=(\|X\|_{F}^{2}+\|Y\|_{F}^{2}+1)^{2}-\sum_{i,k}\log(X_{ik})-\sum_{k,j}\log(Y_{kj})$ and some $\ell,\kappa>0$. The second formulation yields the problem
$\min_{X\in\mathbb{R}_{++}^{m\times r},Y\in\mathbb{R}_{++}^{r\times n}}f_{\text{KL}}(X,Y)=\tfrac{1}{mn}\left(\sum_{i=1}^{m}\sum_{j=1}^{n}D\left(Z_{ij}\Big\|\sum_{k=1}^{r}X_{ik}Y_{kj}\right)\right),$ (3)
where the KL divergence between $x$ and $y$ is defined as $D(x\|y)=x\log(x/y)-x+y$.
@@
To prove that $f_{\text{KL}}(X,Y)$ is $\ell F_{\text{KL}}$-based $\kappa$-self-concordant for some convex function $F_{\text{KL}}$ and for some $\ell,\kappa>0$ on $\mathbb{R}_{++}^{m\times r}\times\mathbb{R}_{++}^{r\times n}$, let $g(x,y)=-\log(x^{\top}y)$ and $G(x,y)=-\sum_{i=1}^{r}\log(x_{i}y_{i})$. By Proposition 2.4, it suffices to prove that $g$ is $\tau G$-based $1$-self-concordant for some $\tau>3$. Fixing $(x,y)\in\mathbb{R}_{++}^{r}\times\mathbb{R}_{++}^{r}$ and $h=(h_{x},h_{y})\in\mathbb{R}^{r}\times\mathbb{R}^{r}$, we define
$s=x^{\top}y,\quad L=h_{x}^{\top}y+h_{y}^{\top}x,\quad u_{i}=\tfrac{|(h_{x})_{i}|}{x_{i}},\quad v_{i}=\tfrac{|(h_{y})_{i}|}{y_{i}},\quad w_{i}=\tfrac{x_{i}y_{i}}{s},\quad S=\textstyle\sum_{i=1}^{r}(u_{i}^{2}+v_{i}^{2}).$
@@
$\tfrac{|L|}{s}\leq\textstyle\sum_{i=1}^{r}w_{i}(|u_{i}|+|v_{i}|)\leq\sqrt{2\left(\textstyle\sum_{i=1}^{r}w_{i}(u_{i}^{2}+v_{i}^{2})\right)}\leq\sqrt{2S},\quad\tfrac{|h_{x}^{\top}h_{y}|}{s}\leq\textstyle\sum_{i=1}^{r}\tfrac{|(h_{x})_{i}||(h_{y})_{i}|}{x_{i}y_{i}}\leq\tfrac{1}{2}S.$
It follows that
$|\nabla^{2}g(x,y)[h,h]|=\left|\tfrac{L^{2}}{s^{2}}-\tfrac{2h_{x}^{\top}h_{y}}{s}\right|\leq 3S,\quad|\nabla^{3}g(x,y)[h,h,h]|=\left|-2\tfrac{L^{3}}{s^{3}}+\tfrac{6Lh_{x}^{\top}h_{y}}{s^{2}}\right|\leq 7\sqrt{2}S^{3/2}.$
In addition, we obtain that $\nabla^{2}G(x,y)[h,h]=S$ and $|\nabla^{3}G(x,y)[h,h,h]|\leq 2(\sum_{i=1}^{r}(u_{i}^{3}+v_{i}^{3}))\leq 2S^{3/2}$. Thus, we have
$\nabla^{2}(g+\tau G)[h,h]\geq(\tau-3)S,\quad\nabla^{3}(g+\tau G)[h,h,h]\leq(2\tau+7\sqrt{2})S^{3/2}.$
This implies that $|\nabla^{3}(g+\tau G)[h,h,h]|\leq 2(\nabla^{2}(g+\tau G)[h,h])^{3/2}$ for some $\tau>3$. This yields the desired result, i.e., $F_{\text{KL}}(X,Y)=-\frac{1}{mn}(\sum_{i,k}\log(X_{ik})+\sum_{k,j}\log(Y_{kj}))$ and $\ell>3$.
2.2 Descent Guarantees for $F$-based $\kappa$-self-concordant Functions
For $F$-based $\kappa$-self-concordant functions, where $F$ is a convex function, we now establish a lower bound on the decrease that is achievable by an appropriate choice of the step length along a direction of descent. For simplicity, we define two auxiliary functions $\omega,\omega_{\star}:[0,\infty)\to\mathbb{R}\cup\{+\infty\}$ as follows:
$\omega(z)=z-\log(1+z)\quad\textnormal{ and }\quad\omega_{\star}(z)=\begin{cases}-z-\log(1-z)&\textnormal{if }z<1,\\ +\infty&\textnormal{otherwise}.\end{cases}$
This notation is standard in convex self-concordant analysis [58, 72]. In addition, we define
@@ 2.2 Descent Guarantees for $F$-based $\kappa$-self-concordant Functions
Suppose that $f$ is $F$-based $\kappa$-self-concordant and $F$ is convex. Then, for any $x\in\textbf{dom}(f)$ and any descent direction $d\in\mathbb{R}^{n}$ satisfying $\nabla f(x)^{\top}d\leq 0$, defining
$\rho=-\nabla f(x)^{\top}d,\quad\delta=\nabla^{2}f(x)[d,d],\quad\Delta=\nabla^{2}F(x)[d,d],\quad\eta=\tfrac{\rho}{\sqrt{\Delta+\delta}},$
we have, for all $t\in[0,\frac{1}{\kappa\sqrt{\delta+\Delta}})$, that $x+td\in\textbf{dom}(f)$ and
@@
Letting $\bar{t}=\frac{\rho}{\delta+\Delta+\kappa\rho\sqrt{\delta+\Delta}}$, we have $f(x+td)\leq f(x)$ for all $t\in[0,\bar{t}]$ and
$f(x+\bar{t}d)\leq f(x)-\rho\bar{t}+\kappa^{-2}\omega_{\star}(\kappa\bar{t}\sqrt{\delta+\Delta})=f(x)-\kappa^{-2}\omega(\kappa\eta)\leq f(x)-\tfrac{\eta^{2}}{2(1+\Gamma_{f}(x))}.$ (6)
Moreover, if $F$ is not only convex , but is also $\kappa_{F}$-self-concordant, we have
@@
Proof It follows from [58, Theorems 5.1.5 and 5.1.9] that $x+td\in\textbf{dom}(f)$ and
@@
Since $F$ is convex, we have $F(x)+t\nabla F(x)^{\top}d-F(x+td)\leq 0$. Plugging this into the above inequality yields inequality (5). We define
@@
$g^{\prime}(t)=-\rho+\left(\tfrac{t(\delta+\Delta)}{1-\kappa t\sqrt{\delta+\Delta}}\right).$
It is easy to verify from the definition of $\bar{t}$ that $g^{\prime}(t)\leq 0$ for all $t\in[0,\bar{t}]$. Since $g(0)=f(x)$, we have $f(x+td)\leq g(t)\leq f(x)$ for all $t\in[0,\bar{t}]$. Plugging $\bar{t}$ into (5) yields (6). In addition, (6) implies
@@
Thus, $\kappa\eta\leq\Gamma_{f}(x)$. Since $\omega(z)\geq\frac{z^{2}}{2(1+\Gamma_{f}(x))}$ for $0\leq z\leq\Gamma_{f}(x)$, we have $\kappa^{-2}\omega(\kappa\eta)\geq\frac{\eta^{2}}{2(1+\Gamma_{f}(x))}$. It follows from [58, Theorem 5.1.8] that
@@
which implies inequality (7). $\square$
@@ 3 Algorithms
We develop two algorithms, Algorithm 1 (RNM) and Algorithm 2 (ARM), for minimizing an unconstrained function $f$ (i.e., solving problem (1)), where $f$ is $F$-based $\kappa$-self-concordant and bounded below, and present proofs of their convergence. Compared to other regularized methods (e.g., see  [63, 56, 11, 12, 30]), our algorithms assume neither $\textbf{dom}(f)=\mathbb{R}^{n}$ nor a Lipschitz Hessian condition. Throughout the rest of this paper, we make Assumption 3.1 below and impose additional assumptions (e.g., $F$ is self-concordant or quadratic) when necessary.
@@
The function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is bounded below and is $F$-based $\kappa$-self-concordant with respect to a convex function $F:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ for some $\kappa>0$.
We also make use of the following definitions in the rest of this section:
$\rho_{j}=-\nabla f(x_{j})^{\top}d_{j},\quad\delta_{j}=\nabla^{2}f(x_{j})[d_{j},d_{j}],\quad\Delta_{j}=\nabla^{2}F(x_{j})[d_{j},d_{j}],\quad\eta_{j}=\tfrac{\rho_{j}}{\sqrt{\Delta_{j}+\delta_{j}}}.$